OpenStudy (anonymous):

There are 9 boys & 15 girls in a class. Three children are chosen at random . what is the probability that (a) one is a boy and two are girls???

OpenStudy (anonymous):

i am not sure but i think: 2/135

OpenStudy (anonymous):

its 945/2024

OpenStudy (anonymous):

yes@uditkulka is right

OpenStudy (anonymous):

yes i know i am right but what is the sollution i mean explanation can u give me please????

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

see 15+9=24 total =24C3=2024

OpenStudy (anonymous):

and 9C1*15C2=945

OpenStudy (anonymous):

understood

OpenStudy (anonymous):

9/15 = 3/5.....thefore probability is 3/8 of being a boy and 5/8 of being a girl

OpenStudy (anonymous):

whats c1 and C2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

9*1/1*15*14/2=945

OpenStudy (anonymous):

combinations ncr formula

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw